# UseRawInputEncoding

### v5

@@ I Introduction
The density fluctuations in the observed Cosmic Microwave Background ade1 -ade2 requires a nearly scale invariant curvature power spectrum within an early Hubble radius, see brandenberger . The current paradigm to provide these fluctuations, as well as a flat, homogeneous and isotropic universe, is inflation [4]. Inflation can be realised using effective models of canonical quantum scalar fields with a plateau potential in Einstein frame brandenberger -martin . From a similar scalar field plateau potential, one can also obtain nearly invariant curvature perturbations in slowly evolving (contracting or expanding) solutions, piao -lyth . For the inflation solution, conceptual challenges such as super-Planckian problem, multiverse problem aside brandenberger -steinhardt , the isotropic power loss of $5\sim 10\%$ in the small multipole $l<40$ range, suppression of quadruple and dipolar asymmetry ade2 , are notable concerns. Attempts to address these issues usually require pre-”slow-roll inflation” physics such as a constant roll phase starobinsky , a fast roll phase linde , a ”slowly expanding” phase or a ”slowly contracting” phase piao . The slowly contracting phase we focus on in this work is the ”Ekpyrotic” phase in khoury -khoury1 , which requires a bounce to connect the contracting phase to the subsequent expanding phase, whether it be a radiation dominant phase or an inflation phase. A bounce can be non-singular if it occurs at energy scales sufficiently lower than the Planck scale, where the evolution of the classical background is more trustable.
@@
To obtain a bounce mechanism presents a major challenge. After the Ekpyrotic phase, the Einstein frame Hubble constant is $H_{E}<0$ and the scale factor of the universe is in a slowly contracting phase. For an effective bounce to occur, one requires a phase where $\dot{H}_{E}>0$ and after the bounce a subsequent field roll in a non negative potential $V\geq 0$, see xue . In the scalar field kinetic energy dominant epoch, with a shrinking scale factor $a$, the field kinetic energy has the largest power $\epsilon$ in $a^{-2\epsilon}$ amongst cosmological fluids, spatial curvature and static field potential. This slowly contracting phase is shown to be an effective smoothing mechanism kist and addresses the gravitational entropy problem steinhardt2021 . A contracting universe will keep contracting unless some ”additional” conditions exist that leads to $\dot{H}_{E}>0$. One early approach is by allowing, for a short time before the bounce, the kinetic or gradient term to take the wrong sign with its cocomitant violation of Null Energy Conditon (NEC) ljjas . More careful analysis vikman shows that the cosmological perturbations along the transition between dark energy field and phantom field is unstable. Some progress is found in a class of models with Kinetic Gravity Braiding vikman2 and Galileons ljjas1 -ljjas2 , though challenge remains vikman3 . Other notable solution involves higher derivatives operator piao2019 . Attempts to obtain a bounce from canonical scalar fields include using a generic positive curvature matsui and adding string inspired negative energy density term brandenberger2 . In Jordan frame, one can implement a bounce without Null Energy Condition (NEC) violation, li -ljjas3 . Yet these solutions are shown to be dual to inflation in Einstein frame. A Jordan frame model with a conformal coupling involving the Higgs field is also considered in bars -bars2 , in which the Einstein frame scalar field plateau potential is described as a double well potential in Jordan frame. A phase transition is effected when the quartic coupling of the scalar field passes between a positive value region and a negative value region. In this work, we consider a Jordan frame double well potential which goes through a phase transition at conformal coupling.
@@
In this work, by noting that the Einstein frame canonical scalar field with a potential in the form of Eq.(1) admits a corresponding Jordan frame scalar field with a double well potential, we consider that the system is in a broken symmetric phase. As energy density increases we study the scenario that this Jordan frame scalar field goes through a phase transition where the non-minimal coupling bounces off its conformal value with zero mass. We find that $\dot{H}_{E}>0$ appears naturally in the broken symmetric phase leading to a bounce into the symmetric phase. In the next section, we recapitulate the correspondence of Eq.(1) in Jordan frame. We consider the scaling region in the broken symmetric and the symmetric phase, the bounce and the subsequent field roll in the effective potential in the symmetric phase. We consider the value of the cosmological constant in the Einstein frame in the symmetric phase and the scenario of a phase transition back to the broken symmetric phase. We end with a conclusion and discussion.
@@ I.1 The Jordan frame double well potential
We note that from Eq.(19) that the ”Pre-Ekpyrotic” phase of khoury -khoury1 corresponds to a stable $3H_{E}^{2}\sim\lambda$ where the sum of the negative potential and the kinetic term is small comparing to $\lambda$
@@
In Jordan frame, this corresponds to a nearly conformal coupling $K\simeq 6$ and a large $\phi^{2}$ ($\varphi>0$) for some fixed $m^{2}$. As $\varphi$ (and $\phi^{2}$) decreases, when the negative potential $-\frac{1}{2}m^{2}e^{-c\varphi}$ dominates over $\lambda$, one enters the ”Ekpyrotic phase” .
@@
After the Ekpyrotic phase, the scale factor continues to shrink in the Einstein frame (which is also true in Jordan frame for slow varying $\phi>0$). The Hubble constant increases towards the Planck scale. As the canonical kinetic term dominates over the potential with decreasing scale factor, the energy density which scales as $\sim a^{-6}$ can only continue to increase since none of the known cosmological components can slow it down. To avoid reaching the Planck energy density, one way is to consider the existence of a high energy density fixed point below the Planck scale. Scalar-Tensor theories in the form of Eq. (9) are used to describe the universe’s evolution from very high energy densities to low energy densities over a large energy density scales. Conformally invariant fixed point $K=6$ and $m^{2}=0$ provides an important solution. In wetterich2 , for a positive potential
@@
functional renormalisation group flow calculation is done for the parameter $y$ where $\frac{1}{y^{2}}=\frac{m^{2}}{\phi^{2}}$, where $m^{2}$ plays the role of the renormalisation scale $k^{2}$. For $m\rightarrow 0$ much faster than $\phi\rightarrow 0$ ($y\rightarrow\infty$), the system will approach the conformally invariant (infra-red) fixed point described above. This infra-red fixed point can be identified to the Einstein-Hilbert action in Einstein frame. The UV limit is reached by $y\rightarrow 0$, which is the small $\phi$ region where $\phi^{2}\ll m^{2}$, which is the UV fixed point of the asymptotically safe quantum gravity wetterich2 -litim2004 .
With an unimpeded energy density rise, the Hubble radius can shrink to the Planck scale. Alternatively, due to double well potential, before the system reaches the Planck energy density, it could go through a phase transition at the conformal point into the symmetric phase with potential described by Eq.(21). The idea of the scalar inflaton having both a symmetric and a broken symmertic phase is not new. A second order phase transition in quantum gravity around some critical field value $\phi^{2}_{c}$ has also been proposed in falls . Early inflation scenario proposed Guth guth2 assumes a first order phase transition from the symmetric phase to the broken symmetric phase. The New inflation scenario linde2 assumes a continuous phase transition also from the symmetric phase to the broken symmetric phase. In a flat space $R_{J}=0$ scenario, this is the most common mechanism used for phase transition studies luscher . In cosmological evolution, as the energy scale varies across a large range that a scalar field with a given symmetry is expected to go through phase transitions, which could lead to inflation guth2 as well as in the production of cosmological defects kibble .
In this work, we shall follow steinhardt in working in the Einstein frame for $H_{E}$ ($\dot{H}_{E}$) evolution and uses the field ($\varphi$) roll in the Einstein potential to consider the perturbation power spectrum in inflation and contraction, However, we shall use the Jordan frame field ($\phi$) potential to folow its phase transition behaviour.
@@ I.2 Scaling region in the Broken symmetric phase
We note that IF the mass parameter $m^{2}$, $c$ and $\lambda$ are constant, we will recover the usual result $\dot{H}_{E}=-\frac{1}{2}\dot{\varphi}^{2}$ and there is no built-in mechanism to switch off the shrinking scale factor.
@@
Starting with large field speeds and a decreasing mass term $m^{2}/\phi^{2}$, in the scaling region where $\epsilon$ remains nearly constant such that $\dot{\epsilon}\sim 0$. From Eq.( 44) we obtain
@@ I.4 The bounce
We recall the Eq.(34)
$\dot{\rho}=\dot{\epsilon}\frac{d}{d\epsilon}a^{-2\epsilon}=-2\epsilon\rho H_{E}-2\dot{\epsilon}\rho\ln\bigg{(}\frac{a}{a_{p}}\bigg{)}.$ (44)
From the Friedmann equation $3H_{E}^{2}=\frac{1}{2}\dot{\varphi}^{2}-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$, we see that $H_{E}=0$ will be reached in the broken symmetric phase. At $H_{E}=0$, for non-zero field speed $\dot{\varphi}>0$, we could have the equation of state parameter $\epsilon\rightarrow\infty$.
Consider the case $H_{E}=0_{+}$, from Eq.(34), we have
$\frac{\dot{\rho}}{H_{E}}=-2\epsilon\rho\bigg{(}1+\frac{1}{H_{E}}\frac{\dot{\epsilon}}{\epsilon}\ln\bigg{(}\frac{a}{a_{p}}\bigg{)}\bigg{)}>0.$ (45)
That is, the energy density can continue to increase $\dot{\rho}>0$ if $\dot{\epsilon}<0$ and $a>a_{p}$. Here $a_{p}$ plays the role of a ”minimum scale factor”. This is possible due to the fact that the value of $\epsilon$ comes down from a very large value to $\epsilon<3$ as $H_{E}$ passes through $H_{E}=0$. The system can then continue to move into the symmetric phase where the mass term in the potential $V_{E}$ becomes positive. This energy density continues to grow until
@@
where this $H_{E}>0$ system reaches a peak $\rho$ value with $\dot{\rho}=0$ ($H_{E}=\sqrt{\rho/3}$) and $\dot{H}_{E}<0$ in the symmetric phase. The bounce from this point leads naturally to an expanding phase. When the exit potential is $V_{E}<\frac{1}{2}\dot{\varphi}^{2}$, we have $\epsilon>2$ the system enters the radiation epoch directly. If $V_{E}>\dot{\varphi}^{2}$, we have $\epsilon<1$ and the system will go through a short inflationary epoch.
We should point out that in wetterich3 , if the energy density after the bounce is large enough $V_{E}\gg\dot{\varphi}^{2}$, the system possesses a global solution which has a long inflationary phase and eventually crossovers and enters into a late dynamical dark energy epoch.
@@ I.5 After the bounce
For an exit potential $V_{E}$ that is slightly larger than $\rho_{c}$, which is stipulated by KS, after the bounce, the system goes into either a short inflation or the radiation dominant phase. For early bounce universe model, the Einstein frame potential is
@@
with small fixed $m^{2}$. The potential will reduce to $\lambda$ asymptotically when $c\rightarrow 0$ ($K\rightarrow-\infty$) where $\varphi\rightarrow-\infty$, where the energy density $\rho\rightarrow 0$ is implicitly assumed.
Here we can assume explicitly that the $m^{2}$ term has an energy density dependence, so that as $\rho\rightarrow 0$, $m^{2}(\rho)\rightarrow\rho^{l}$ which is faster than $e^{-c\varphi}=\phi^{-2}\rightarrow\rho^{-q}$, such that the term $\frac{1}{2}m^{2}e^{-c\varphi}$ is suppressed in the low energy density limit. $\lambda$ continues to play the role of the cosmological constant.
@@
shows that both terms on the R.H.S. provide friction on the field speed $\dot{\varphi}<0$. The field speed will eventually reach $\dot{\varphi}>0$. We wish to consider whether the system can return to the starting point of Khoury-Steinhardt khoury -khoury1 after a discussion on the cosmological constant.
@@ I.6 Implication for the cosmological constant
Here, we come to a convenient place to consider the issue of cosmological constant value. In Jordan frame, the Einstein frame cosmological constant $\lambda$ becomes the four point coupling constant of a $\phi^{4}$ theory. There are attempts to use the physics of the renormalised $\lambda\phi^{4}$ theory to address the small cosmological constant problem.
Polyakov in polyakov considers the conformally flat universe where $g_{\mu\nu}=\phi^{2}\delta_{\mu\nu}$. In Euclidean time, the Einstein-Hilbert action becomes
@@
In wetterich2 , the authors use the functional renormalisation group to consider the Jordan frame symmetric potential without the $\lambda\phi^{4}$ term. They find that in the infra-red regime and large $\phi$ region, the Einstein frame effective potential $\lambda=V_{E}\sim\frac{m_{g}^{2}}{\phi^{2}}$, where $m_{g}$ is the scalar graviton mass and $\phi=M_{pl}$.
@@
Since $\lambda$ is the measured value at low energies, Hsu and Zee zee postulates that given the Planck mass $M_{p}$ and the mass scale of the observable universe $M_{U}=2\times 10^{-33}eV$, which matches the smallest upper bound of the massive graviton, de Rham et al. deRham . The vacuum energy density takes the value
@@
Our approach is different. Here, the Jordan frame potential is in the symmetric phase at low energy density, the renormalised scalar gravitation mass is small and approaching zero. As discussed in luscher , if we assume that the $\lambda\phi^{4}$ system is sitting in the scaling region of a phase transition to a broken symmetric phase, the renormalised $\lambda$ takes the value the coupling constant which matches its counterpart
@@
in the scaling region of broken symmetric phase luscher . We take the scalar graviton mass to be close to $10^{-33}eV$ from different constraints in deRham , so that we have

### v6

@@ I Introduction
The density fluctuations in the observed Cosmic Microwave Background ade1 -ade2 requires a nearly scale invariant curvature power spectrum within an early Hubble radius, see brandenberger . The current paradigm to provide these fluctuations, as well as a flat, homogeneous and isotropic universe, is inflation [4]. Inflation can be realised using effective models of canonical quantum scalar fields with a plateau potential in Einstein frame brandenberger -martin . From a similar scalar field plateau potential, one can also obtain nearly invariant curvature perturbations in slowly evolving (contracting or expanding) solutions, piao -lyth . For the inflation solution, conceptual challenges such as super-Planckian problem, multiverse problem aside brandenberger -steinhardt , the isotropic power loss of $5\sim 10\%$ in the small multipole $l<40$ range, suppression of quadruple and dipolar asymmetry ade2 , are notable concerns. Some attempts to address these issues involve pre-”slow-roll inflation” physics such as a constant roll phase starobinsky , a fast roll phase linde , a ”slowly expanding” phase or a ”slowly contracting” phase piao . The slowly contracting phase we focus on in this work is the ”Ekpyrotic” phase in khoury -khoury1 , which requires a bounce to connect the contracting phase to the subsequent expanding phase, whether it be a radiation dominant phase or an inflation phase. A bounce can be non-singular if it occurs at energy scales sufficiently lower than the Planck scale.
@@
To obtain a bounce mechanism presents a major challenge. After the Ekpyrotic phase, the Einstein frame Hubble constant is $H_{E}<0$ (and $\dot{H}_{E}<0$) and the scale factor of the universe is in a slowly contracting phase. For an effective bounce to occur, one requires a phase where $\dot{H}_{E}>0$ and after the bounce a subsequent field roll in a non negative potential $V\geq 0$, see xue . In the scalar field kinetic energy dominant epoch, with a shrinking scale factor $a$, the field kinetic energy has the largest power $\epsilon$ in $a^{-2\epsilon}$ amongst cosmological fluids, spatial curvature and static field potential. This slowly contracting phase is shown to be an effective smoothing mechanism kist and addresses the gravitational entropy problem steinhardt2021 . A contracting universe will keep contracting unless some ”additional” conditions exist that leads to $\dot{H}_{E}>0$. One early approach is by allowing, for a short time before the bounce, the kinetic or gradient term to take the wrong sign with its cocomitant violation of Null Energy Conditon (NEC) ljjas . More careful analysis vikman shows that the cosmological perturbations along the transition between dark energy field and phantom field is unstable. Some progress is found in a class of models with Kinetic Gravity Braiding vikman2 and Galileons ljjas1 -ljjas2 , though challenge remains vikman3 . Other notable solution involves higher derivatives operator piao2019 . Attempts to obtain a bounce from canonical scalar fields include using a generic positive curvature matsui and adding string inspired negative energy density term brandenberger2 . In Jordan frame, one can implement a bounce without Null Energy Condition (NEC) violation, li -ljjas3 . Yet these solutions are shown to be dual to inflation in Einstein frame. A Jordan frame model with a conformal coupling involving the Higgs field is also considered in bars -bars2 , in which the Einstein frame scalar field plateau potential is described as a double well potential in Jordan frame. A phase transition is effected when the quartic coupling of the scalar field passes between a positive value region and a negative value region. In this work, we consider a Jordan frame double well potential which goes through a phase transition at conformal coupling.
@@
In this work, by noting that the Einstein frame canonical scalar field with a potential in the form of Eq.(1) admits a corresponding Jordan frame scalar field with a double well potential, we consider that the system is in a broken symmetric phase. As energy density increases we study the scenario that this Jordan frame scalar field goes through a phase transition where the non-minimal coupling bounces off its conformal value with zero mass. We find that $\dot{H}_{E}>0$ appears naturally in the broken symmetric phase leading to a bounce into the symmetric phase. In the next section, we recapitulate the correspondence of Eq.(1) in Jordan frame. We consider the scaling region in the broken symmetric and the symmetric phase, the bounce and the subsequent field roll in the effective potential in the symmetric phase. We consider the value of the cosmological constant in the Einstein frame in the symmetric phase and briefly consider the scenario of a phase transition back to the broken symmetric phase. We end with a conclusion and discussion.
@@ I.1 The Jordan frame double well potential
We note that from Eq.(19) that the ”Pre-Ekpyrotic” phase of khoury -khoury1 corresponds to a stable $3H_{E}^{2}\sim\lambda$ where the sum of the negative (mass) potential term and the kinetic term is small comparing to $\lambda$
@@
In Jordan frame, this corresponds to a nearly conformal coupling $K\simeq 6$ and a large $\phi^{2}$ ($\varphi>0$) for some small $m^{2}$. As $\varphi$ (and $\phi^{2}$) decreases, when the negative potential $-\frac{1}{2}m^{2}e^{-c\varphi}$ dominates over $\lambda$, one enters the ”Ekpyrotic phase” .
@@
After the Ekpyrotic phase, the scale factor continues to shrink in the Einstein frame (which is also true in Jordan frame for slow varying $\phi>0$). The Hubble constant increases towards the Planck scale. As the canonical kinetic term dominates over the potential with decreasing scale factor, the energy density which scales as $\sim a^{-6}$ can only continue to increase since none of the known cosmological components can slow it down.
To avoid reaching the Planck energy density, one way is to consider the existence of a high energy density fixed point below the Planck scale. Scalar-Tensor theories in the form of Eq. (9) are used to describe the universe’s evolution from very high energy densities to low energy densities over a large energy density scales. Conformally invariant fixed point $K=6$ and $m^{2}=0$ provides an important solution. In wetterich2 , for a positive potential
@@
functional renormalisation group flow calculation is done for the parameter $y$ where $\frac{1}{y^{2}}=\frac{m^{2}}{\phi^{2}}$, where $m^{2}$ plays the role of the renormalisation energy scale $k^{2}$, here the direct proportionality of $m^{2}$ to $k^{2}$ is assumed. For $m\rightarrow 0$ much faster than $\phi\rightarrow 0$ ($y\rightarrow\infty$), the system will approach the conformally invariant (infra-red) fixed point described above. This infra-red fixed point can be identified to the Einstein-Hilbert action in Einstein frame. The UV limit is reached by $y\rightarrow 0$, which is the small $\phi$ region where $\phi^{2}\ll m^{2}$, which is the UV fixed point of the asymptotically safe quantum gravity wetterich2 -litim2004 .
To avoid the Hubble radius shrinks to the Planck scale, another way is to consider, due to double well potential, before the system reaches the Planck energy density, it could go through a phase transition at the conformal fixed point (where $m^{2}=0$, $K=6$, $H_{E}=0$) into the symmetric phase with potential described by Eq.(21). The idea of the scalar inflaton having both a symmetric and a broken symmertic phase is not new. A second order phase transition in quantum gravity around some critical field value $\phi^{2}_{c}$ has also been proposed in falls . Early inflation scenario proposed Guth guth2 assumes a first order phase transition from the symmetric phase to the broken symmetric phase. The New inflation scenario linde2 assumes a continuous phase transition also from the symmetric phase to the broken symmetric phase. In a flat space $R_{J}=0$ scenario, this is the most common mechanism used for phase transition studies luscher . In cosmological evolution, as the energy scale varies across a large range that a scalar field with a given symmetry is expected to go through phase transitions, which could lead to inflation guth2 as well as in the production of cosmological defects kibble .
In this work, we shall follow steinhardt in working in the Einstein frame for $H_{E}$ ($\dot{H}_{E}$) evolution and uses the field ($\varphi$) roll in the Einstein potential to consider the perturbation power spectrum in inflation and contraction. However, we shall use the Jordan frame field ($\phi$) potential to follow its phase transition behaviour.
@@ I.2 Scaling region in the Broken symmetric phase
We note that IF the mass parameter $m^{2}$, $c$ and $\lambda$ are constant, we will recover the usual result $\dot{H}_{E}=-\frac{1}{2}\dot{\varphi}^{2}$ and there is no built-in mechanism to switch off the shrinking of the scale factor.
@@
Starting with large field speeds and a decreasing mass term $m^{2}/\phi^{2}$, in the scaling region where $\epsilon$ remains nearly constant such that $\dot{\epsilon}\sim 0$. From Eq.( 28) we obtain
@@ I.4 The bounce
From the Friedmann equation $3H_{E}^{2}=\frac{1}{2}\dot{\varphi}^{2}-\frac{1}{2}m^{2}e^{-c\varphi}+\lambda$, we see that for non-zero field speed $\dot{\varphi}>0$, $H_{E}=0$ can in principle be reached in the broken symmetric phase, where the equation of state parameter $\epsilon\rightarrow\infty$. However, the mass term $-m^{2}/\phi^{2}$ freezes and changes sign, the $H_{E}=0$ will not be reached but $H_{E}<0$ simply changes to $H_{E}>0$ and implements the bounce. Within a finite size Hubble radius, the asymptotic limit $\lambda$ will be finite.
We consider the scenario that $H_{E}=0$ can be reached and the mass term in Eq.(35) can move continuously from negative value towards positive value. From Eq.(34), we have initially $H_{E}=0_{+}$
$\dot{\rho}=-2\epsilon\rho\bigg{(}H_{E}+\frac{\dot{\epsilon}}{\epsilon}\ln\bigg{(}\frac{a}{a_{p}}\bigg{)}\bigg{)}>0.$ (44)
Here, the energy density can continue to increase $\dot{\rho}>0$ if $\dot{\epsilon}<0$ and $a>a_{p}$. Here $a_{p}$ plays the role of a ”minimum scale factor”. $\dot{\epsilon}<0$ is possible since the mass term increases in Eq.(35) will take the correspoding value of $\epsilon$ to come down from infinity to $\epsilon<3$ as $H_{E}$ passes through $H_{E}=0$ to $H_{E}>0$. The system reaches the symmetric phase where the mass term in the potential $V_{E}$ takes positive value. This energy density continues to grow until
@@
where this $H_{E}>0$ system reaches a peak $\rho$ value with $\dot{\rho}=0$ ($H_{E}\geq\sqrt{\rho/3}$) and $\dot{H}_{E}<0$ in the symmetric phase. The bounce from this point leads naturally to an expanding phase. When $\dot{\rho}=0$, if the potential takes the value $V_{E}<\frac{1}{2}\dot{\varphi}^{2}$, we have $\epsilon>2$ and the system enters the radiation epoch directly. If $V_{E}>\dot{\varphi}^{2}$, we have $\epsilon<1$ and the system will go through a short inflationary epoch.
We should point out from the results in Wetterich wetterich3 that if the energy density after the bounce is large enough such that $V_{E}\gg\dot{\varphi}^{2}$, the system possesses a global solution which initially has a long inflationary phase and eventually crossovers and enters into a late dynamical dark energy epoch.
@@ I.5 After the bounce
After $\dot{\rho}=0$, for an potential $V_{E}$ that is slightly larger than $\rho_{c}$, which is stipulated by KS, after the bounce, the system goes into either a short inflation or the radiation dominant phase. For bounce universe models considered in the literature (ref), the Einstein frame potential is
@@
where small $m^{2}$ remains fixed from the beginning. The potential will reduce to $\lambda$ asymptotically when $c$ moves away from infinity and $\varphi\rightarrow-\infty$, where the energy density $\rho\rightarrow 0$ is assumed.
Here we can assume explicitly that the $m^{2}$ term also has an energy density dependence, so that as $\rho\rightarrow 0$, we have also $m^{2}(\rho)\rightarrow\rho^{l}$ ($l>0$) in addition to $e^{-c\varphi}=\phi^{-2}\rightarrow\rho^{q}$, such that the term $\frac{1}{2}m^{2}e^{-c\varphi}$ reduces more rapidly in the low energy density limit. $\lambda$ continues to play the role of the cosmological constant.
@@
shows that both terms on the R.H.S. provide friction on the field speed $\dot{\varphi}<0$. The negative field speed $\dot{\varphi}<0$ will eventually reach $\dot{\varphi}>0$. We wish to consider whether the system can return to the starting point of Khoury-Steinhardt khoury -khoury1 after a quick discussion on the cosmological constant.
@@ I.6 Implication for the cosmological constant
Here, we come to a convenient place to consider the issue of cosmological constant value. In Jordan frame, the Einstein frame cosmological constant $\lambda$ becomes the four point coupling constant of a $\phi^{4}$ theory.
There are attempts to use the physics of the renormalised $\lambda\phi^{4}$ theory to address the small cosmological constant problem. Polyakov in polyakov considers the conformally flat universe where $g_{\mu\nu}=\phi^{2}\delta_{\mu\nu}$. In Euclidean time, the Einstein-Hilbert action becomes
@@
In wetterich2 , the authors use the functional renormalisation group to consider the Jordan frame symmetric potential without the $\lambda\phi^{4}$ term. They find that in the infra-red regime and large $\phi$ region, the Einstein frame cosmological constant is the effective potential $V_{E}\sim\frac{m_{g}^{2}}{\phi^{2}}$, where $m_{g}$ is the scalar graviton mass and $\phi=M_{pl}$.
@@
Since $\lambda$ is the measured value at low energies, Hsu and Zee zee postulates that given the Planck mass $M_{pl}$ and the mass scale of the observable universe $M_{U}=2\times 10^{-33}eV$, which matches the smallest upper bound of the massive graviton, de Rham et al. deRham . The vacuum energy density takes the value
@@
Our approach is slightly different. Here, the Jordan frame potential is in the symmetric phase at low energy density, if the renormalised scalar graviton mass term $m_{g}^{2}/\phi^{2}$ is much larger than $\lambda$, the effective cosmological constant term is
$\lambda=V(\varphi)\sim\frac{1}{2}\frac{m^{2}}{\phi^{2}}.$ (54)
If the renormalised mass term is much smaller than $\lambda$, as discussed in luscher , if we assume that the $\lambda\phi^{4}$ system is sitting in the scaling region of a phase transition to a broken symmetric phase, the renormalised coupling constant $\lambda$ takes on the same value as its counterpart in the scaling region of broken symmetric phase luscher , such that, in terms of the definition in luscher
@@
We take the scalar graviton mass to be close to $10^{-33}eV$ in deRham , so that we have an estimate of the cosmological constant as